OpenStudy (anonymous):

36y^2-84y-147 factor

OpenStudy (anonymous):

First of all u can factor this as - 3(12y^2-28y-49) -49x12 = -588. and even -42 x 14 = -588. => 3 (12^y2 - 42 y + 14 y - 49) solve accordingly :)

OpenStudy (anonymous):

36y^2-84y-147 ={(6y)^2-2*6y*7+7^2-7^2}-147 =(6y-7)^2-49-147 =(6y-7)^2-196 =(6y-7)^2-14^2 =(6y-7+14)(6y-7-21) =(6y+7)(6y-21)

OpenStudy (anonymous):

=(6y+7)3(2y-7) =3(6y+7)(2y-7)